OpenStudy (anonymous):

The rational function has a y-intercept of 7. What is the equation for this function?

OpenStudy (tkhunny):

Ummm... There are infinitely many. Can you narrow it down a bit?

OpenStudy (anonymous):

this is the graph. ^

OpenStudy (whpalmer4):

Do you recognize that the function is shifted by -2, 5 from the origin? Do you know the form of the function if centered on the origin?

OpenStudy (tkhunny):

It looks like this: \(y = \dfrac{ax+b}{cx+d}\) y-intercept is 7, substitute x = 0 \(7 = \dfrac{b}{d} \implies b = 7d\) Now, it looks like this: \(y = \dfrac{ax+7d}{cx+d}\) Vertical Asymptote at x = -2, the denominator must be zero at x = 2 \(c(-2) + d = 0 \implies d = 2c\) Now, it looks like this: \(y = \dfrac{ax+7(2c)}{cx+2c} = \dfrac{ax+14c}{c(x+2)}\) Horizontal Asymptote at y = 5. This one is a little trickier. \(a = 5c\) Now, it looks like this: \(y = \dfrac{(5c)x+14c}{c(x+2)} = \dfrac{c(5x + 14)}{c(x+2)} = \dfrac{5x+14}{x+2}\). I think it's done!

OpenStudy (anonymous):

Thats really confusing. . . .

OpenStudy (tkhunny):

All at once, I agree. One step at a time and it's a beautiful thing. Think through each step. Don't try to do it all at the same time.

OpenStudy (tkhunny):

* Whoops! The denominator must be zero at x = -2.